# Title of Your Paper

### v1

@@ 4.1 Fairness Incompatibility and Trade-off Analysis
Table 4.1.1: Assessing Equity: A Comparative Analysis of Fairness Metrics Between Non-Hispanic White and Black Applicants.
@@
Table 4.1.2: Evaluating Inclusivity: Disparity Measures in Fairness Metrics Between Non-Hispanic White Applicants and People of Color.
@@
4.2 Model Accuracy under Race-Aware vs. Race-Blind Conditions
4.2.1 Algorithmic Accuracy Comparisons
Figure 2: Algorithm Accuracy without Demographic Features of Race, Sex, and Minority Population of the Census Tract Area
@@ 4.2.1 Algorithmic Accuracy Comparisons
The random forest model has the highest median accuracy above 67%, indicating it is generally the most accurate among the algorithms presented (Figure 2). Logistical regression and KNN models also reflect relatively high and consistent accuracy. In contrast, the CART model exhibits a wider range of accuracy values, suggesting it may be more sensitive to the specific data or parameters used. The NB model demonstrates the lowest median accuracy among the five models tested.
@@
Figure 3: Change in algorithm accuracy after inclusion of race
@@
The logistical model exhibits minimal variation, with most of the data clustered around zero. This implies that incorporating race has little to no effect on fairness or model performance, making it a neutral choice in terms of impact. Similarly, KNN and NB also demonstrate minimal variation, making their stability comparable to the logistic regression model in terms of consistency and reliability. This suggests that including race has a neutral to minimal impact on fairness for these three models, making them consistent options for scenarios where stability is key.
@@
Figure 4: Trade-off: financial inclusion vs. negative impact on minorities— with and without Race
@@ 5 Discussion
Lee and Floridi (2021)’s technique uses relative trade-offs between objectives as a marker for fairness. This approach emphasizes flexibility for each individual context, allowing distinct variables to be considered in relation, and applies a relative evaluation scheme by comparing different algorithms. Based on a decision-maker’s specific goals and context, a different algorithm may be suggested, providing the best balance between different objectives and thus may represent the most “fair” approach for that use case. There is no single ”one-size-fits-all” method that guarantees optimal fairness. Instead, decision-makers must balance objectives according to their unique circumstances, and consider what information is required to achieve fairer outcomes in their models (e.g., inclusion or exclusion of race depending on model family). A larger loan can generate higher revenue for a company, but loan rejections also carry the risk of financial loss. A higher loan size, in turn, can lead to a greater rejection rate, as seen in the comparison between models e.g., random forest (RF) and Naïve Bayes (NB) without race. In the lending context, the trade-off between financial inclusion of minority borrowers and the accuracy of predicting loan defaults requires careful calibration, and a moderate approach may be preferred. For cases similar to the one described in this study, this may mean use of CART, random forest, or KNN models, for example.
@@
The assumption that an inherent trade-off exists between reducing disparity and maintaining accuracy has also been challenged by recent computer science research. In this context, disparate impact refers to the disproportionate negative effects a policy or model may have on disadvantaged groups, while accuracy reflects the model’s effectiveness in achieving its predictive goals, such as assessing loan default risk. Model multiplicity demonstrates that multiple models can achieve similar predictive accuracy while exhibiting vastly different levels of disparate impact (Black et al., 2022). This means that companies are not necessarily forced to choose between fairness, performance, and accuracy—they can be optimized in conjunction if properly contextualized fairness considerations are integrated into model development from the outset. Thus, rather than settling for models that perpetuate bias, institutions should actively search for less discriminatory algorithms that maintain predictive accuracy while reducing disparities.
@@
5.1 Future Directions and Recommendations
@@
While the financial benefits of adopting AI in the financial services industry are well-documented, including increased efficiency and broadening credit, mortgage, and financial access to traditionally excluded customers, a recent US Treasury (2024) survey report highlighted that lenders and insurers are still reluctant to adopt AI in their decision-making processes. The responsibility associated with creating and deploying public-facing automated decision-making tools for resource allocation or opportunity access is a heavy one for public and private sector entities to hold. AI is a technology that in many ways is still in its infancy—developing sensible internal risk management infrastructure in model procurement, development, and deployment processes are outstanding considerations before mass adoption.
@@
These concerns ultimately revolve around questions of liability—where to assign legal responsibility for the decisions of automated systems, and how to successfully and precisely mitigate unintended harms before they occur. The current legal regime applicable to automated systems consists of existing civil rights statutes such as Title VII, the Equal Credit Opportunity Act (ECOA), and the Fair Housing Act. Also directly applicable have been interpretations of algorithmic harm through the lens of disparate impact theory, as seen in Title VI ((Black et al., 2024b)). In the past two years, LDAs in particular have received attention from AI practitioners and federal bodies as a paradigm to achieve compliance with existing civil rights law. CFPB’s annual Fair Lending report (Consumer Financial Protection Bureau, 2024, p.36) made what responsibilities that entails clear for lenders and insurers, ”Robust fair lending testing of models should include regular testing for disparate treatment and disparate impact, including searches for and implementation of less discriminatory algorithms using manual or automated techniques.”
@@
This study has made the case for one high-impact evaluation framework in the statistical modeling stage of AI model development, a re-contextualized fairness framework that is capable of being easily integrated into the existing algorithmic development process pre-deployment for public interacting models. Previous corporate attempts (e.g., Upstart) to include fairness interventions with LDAs have highlighted the critical nature of not only having solutions that are impactful in improving outcomes across protected classes, but making sure they are time efficient, cost-effective, and accessible to organizations of all sizes (Relman Colfax, PLLC, 2024). Conducting a horizontal LDA search is one such lightweight fairness intervention that can be taken to guide LDA searches for resource-constrained institutions, or as a first step before more computational intensive approaches—such as vertical LDA searches—for organizations wanting to commit to a more comprehensive search process.
@@
In hopes of trying to amplify best practices and encourage broader adoption of responsible AI development methodologies, this paper also offers four recommendations for public and private sector entities trying to develop their risk management frameworks and LDA searches within the existing AI legal landscape.
5.2 Recommendation 1: Develop an LDA Forward Development Process
@@
Any serious movement towards transitioning to automated decision-making systems in the public and private sectors merits utilizing a rigorous, justifiable, and transparent process in the early model development stage. The foundation of this is predicated on a structured and well-documented LDA search. What this looks like might vary based on institutional capacity to establish what is “reasonable” for their particular case. For some, an initial horizontal LDA search, surveying across models for fairness gains, may be sufficient, while others may want to consider more comprehensive approaches as used in vertical LDA searches, e.g., tuning, or some combination of both. Regardless, a search should be undertaken as the current disparate impact legal regime expects that should precautions not be taken or deemed insufficient, an organization could then expect to bear costs associated with re-evaluating, re-developing, and replacing the product or service with an LDA (Black et al., 2024b). It is preferable then that rather than looking at fairness assessments as an isolated step in the model development cycle, organizations should aim to have interoperable internal systems in place that can responsibly steward and deploy data across the model lifecycle.
@@
5.3 Recommendation 2: Consider a Developer-first Approach to Documentation and Compliance
State-based AI legislation is demonstrating an emerging interest in AI accountability as primarily a developer-oriented responsibility (Colorado General Assembly, 2024). With model multiplicity being a widely accepted theory in the AI community, there is almost always a less discriminatory algorithm should one look for it; thus the law is increasingly resting the burden of model harm in the development process if efforts to circumvent potential harm are not proven. Against this context, organizations should consider adjusting their risk management and compliance strategies to focus on the role of their engineering team in the product development process.
@@ 5.3 Recommendation 2: Consider a Developer-first Approach to Documentation and Compliance
5.4 Recommendation 3: Ensure Adequate Government Support for LDA search & Development
To ensure that financial institutions conduct robust searches for LDAs, fairness obligations must be accompanied by clear guidance, concrete examples, and best-practice principles (Chien et al., 2024). Policymakers should provide comprehensive guidelines in key areas, including appropriate debiasing techniques, recommendations for the proper depth of LDA searches, considerations for LDA viability, and suitable fairness metrics in a variety of different contexts (Chien et al., 2024).
@@ 5.4 Recommendation 3: Ensure Adequate Government Support for LDA search & Development
5.5 Recommendation 4: Industry peers should benchmark algorithms and share best practices
@@
Each of these recommendations are independent, yet complementary steps for the at-scale adoption of AI accountability in practice. They assist in translating the normative values responsible AI encourages into systems and processes, establishing ubiquitous, standardized AI fairness across the board. As outlined, developers, institutions, and governments each have role-specific responsibilities to develop systemic and structured expectations on the process, documentation, and standards required to achieve this goal. Auditing and compliance work, as a result, can become less opaque, and the use and development of AI systems can be demystified for increased process efficiency, liability, and public trust cultivation.
6 Conclusion
7 Appendix
Appendix A Data Visualizations
A.1 HMDA 2021 Home Ownership Race Distribution for Approved and Denied Loans
Figure A. 1:
A.2 HMDA 2021 Home Ownership Race Distribution for Average Loan Amount
Figure A. 2:
A.3 HMDA 2021 Home Ownership Race Distribution for Average Interest Rates
Figure A. 3:
Appendix B Model Development
B.1 Features used for Model Development

### v2

@@ 4.1. Fairness Incompatibility and Trade-off Analysis
Table 4.1.1. Assessing Equity: A Comparative Analysis of Fairness Metrics Between Non-Hispanic White and Black Applicants.
@@
Table 4.1.2. Assessing Equity: Disparity Measures in Fairness Metrics Between Non-Hispanic White Applicants and People of Color.
@@
4.2. Model Accuracy under Race-Aware vs. Race-Blind Conditions
4.2.1. Algorithmic Accuracy Comparisons
Figure 2. Algorithm Accuracy without Demographic Features of Race, Sex, and Minority Population of the Census Tract Area
@@ 4.2.1. Algorithmic Accuracy Comparisons
The random forest model has the highest median accuracy above 67%, indicating it is generally the most accurate among the algorithms presented (Figure 2). Logistic regression and KNN models also reflect relatively high and consistent accuracy. In contrast, the CART model exhibits a wider range of accuracy values, suggesting it may be more sensitive to the specific data or parameters used. The NB model demonstrates the lowest median accuracy among the five models tested.
@@
Figure 3. Change in Algorithm Accuracy after Inclusion of Race
@@
The logistic model exhibits minimal variation, with most of the data clustered around zero. This implies that incorporating race has little to no effect on fairness or model performance, making it a neutral choice in terms of impact. Similarly, KNN and NB also demonstrate minimal variation, making their stability comparable to the logistic regression model in terms of consistency and reliability. This suggests that including race has a neutral to minimal impact on fairness for these three models, making them consistent options for scenarios where stability is key.
@@
Figure 4. Financial Inclusion vs. Negative Impact on Minorities— with and without Race
@@ 5. Discussion
lee21’s technique uses relative trade-offs between objectives as a marker for fairness. This approach emphasizes flexibility for each individual context, allowing distinct variables to be considered in relation, and applies a relative evaluation scheme by comparing different algorithms. Based on a decision-maker’s specific goals and context, a different algorithm may be suggested, providing the best balance between different objectives and thus may represent the most “fair” approach for that use case. There is no single ”one-size-fits-all” method that guarantees optimal fairness. Instead, decision-makers must balance objectives according to their unique circumstances, and consider what information is required to achieve fairer outcomes in their models (e.g., inclusion or exclusion of race depending on model family). A larger loan can generate higher revenue for a company, but loan rejections also carry the risk of financial loss. A higher loan size, in turn, can lead to a greater rejection rate, as seen in the comparison between models e.g., random forest (RF) and Naïve Bayes (NB) without race. In the lending context, the trade-off between financial inclusion of minority borrowers and the accuracy of predicting loan defaults requires careful calibration, and a moderate approach may be preferred. For cases similar to the one described in this study, this may mean use of CART, random forest, or KNN models, for example.
@@
The assumption that an inherent trade-off exists between reducing disparity and maintaining accuracy has also been challenged by recent computer science research. In this context, disparate impact refers to the disproportionate negative effects a policy or model may have on disadvantaged groups, while accuracy reflects the model’s effectiveness in achieving its predictive goals, such as assessing loan default risk. Model multiplicity demonstrates that multiple models can achieve similar predictive accuracy while exhibiting vastly different levels of disparate impact (black2022). This means that companies are not necessarily forced to choose between fairness, performance, and accuracy—they can be optimized in conjunction if properly contextualized fairness considerations are integrated into model development from the outset. Thus, rather than settling for models that perpetuate bias, institutions should actively search for less discriminatory algorithms that maintain predictive accuracy while reducing disparities.
@@
5.1. Future Directions and Recommendations
@@
The financial benefits of adopting AI in the financial services industry are well-documented, including increased efficiency and broadening credit, mortgage, and financial access to traditionally excluded customers. A recent US schaffer2024 survey report, however, highlighted that lenders and insurers are still reluctant to adopt AI in their decision-making processes. The responsibility associated with creating and deploying public-facing automated decision-making tools for resource allocation or opportunity access is a heavy one for public and private sector entities to hold. AI is a technology that in many ways is still in its infancy, and developing sensible risk management infrastructure in model procurement, development, and deployment processes are outstanding considerations before mass adoption.
@@
These concerns ultimately revolve around questions of liability—where to assign legal responsibility for the decisions of automated systems, and how to successfully and precisely mitigate unintended harms before they occur. The current legal regime applicable to automated systems consists of existing civil rights statutes such as Title VII, the Equal Credit Opportunity Act (ECOA), and the Fair Housing Act. Also directly applicable have been interpretations of algorithmic harm through the lens of disparate impact theory, as seen in Title VI ((black2024)). In the past two years, LDAs in particular have received attention from AI practitioners and federal bodies as a paradigm to achieve compliance with existing civil rights law. CFPB’s annual Fair Lending report (cfpb2024, p.36) made what responsibilities that entails clear for lenders and insurers, ”Robust fair lending testing of models should include regular testing for disparate treatment and disparate impact, including searches for and implementation of less discriminatory algorithms using manual or automated techniques.”
@@
This study has made the case for one high-impact evaluation framework in the statistical modeling stage of AI model development, a re-contextualized fairness framework that is capable of being easily integrated into the existing algorithmic development process pre-deployment for public interacting models. Previous corporate attempts (e.g., Upstart) to include fairness interventions with LDAs have highlighted the critical nature of not only having solutions that are impactful in improving outcomes across protected classes, but making sure they are time efficient, cost-effective, and accessible to organizations of all sizes (colfax2024). Conducting a horizontal LDA search is one such lightweight fairness intervention that can be taken to guide LDA searches for resource-constrained institutions, or as a first step before more computational intensive approaches—such as vertical LDA searches—for organizations wanting to commit to a more comprehensive search process. For relevant policy and development recommendations, see Appendix B
6. Conclusion
Appendix A A Review of Machine Learning Algorithms
A.0.1. Logistic regression (LR)
A.0.2. K-nearest neighbors classification (KNN)
The K-nearest neighbors (KNN) algorithm is a non-parametric, supervised machine learning approach utilized for classification and regression tasks. In the KNN algorithm, data points are assigned to a class based on the majority class of their ’K’ nearest neighbors in the feature space. The ’K’ in KNN signifies the number of neighbors considered for the classification or regression decision. In the context of our focus on classification tasks, the algorithm assigns a class label to a data point through a majority vote from its nearest neighbors. For regression tasks, KNN computes the average value of the K nearest neighbors to predict a continuous outcome. KNN is versatile and does not assume any underlying data distribution, making it suitable for various types of datasets. However, its performance can be sensitive to the choice of the distance metric and the value of K, which may become computationally expensive for large datasets. Additionally, it suffers from the curse of dimensionality if the dimensions of the feature space are large, as there will generally be very few training observations close to the point of interest. For a recent review and in-depth analysis of KNN, we refer interested readers to syriopoulos2023k. For this study, we use K=5 as specified in lee21.
A.0.3. Classification and regression tree (CART)
Classification and regression trees (CART) introduced by breiman1984classification constitutes a nonparametric supervised learning approach that builds tree-like models for both categorical (classification) and continuous (regression) prediction tasks. The algorithm recursively partitions the feature space into distinct regions, assigning a predictive model (either class labels for classification or continuous values for regression) to each resulting segment. CART employs binary splitting at each node, selecting the optimal feature and threshold based on criteria such as Gini impurity for classification or mean squared error for regression. The resulting tree structure provides interpretable decision paths, helping to understand complex relationships within the data. While CART inherently handles non-linear relationships and interactions, it may be prone to overfitting, prompting the adoption of techniques like pruning and ensemble methods for regularization. However, it is an unstable algorithm, particularly when dealing with noisy data or in the presence of outliers.
A.0.4. Gaussian Naïve Bayes (NB)
Gaussian Naïve Bayes (NB) is a probabilistic generative model designed for classification tasks, known for its simplicity and efficiency. It is derived from Bayes’ theorem and assumes that features are conditionally independent given the class label, thereby simplifying the computation of probabilities. In the context of continuous features, Gaussian NB models the likelihood of each feature’s distribution as Gaussian, facilitating the estimation of class probabilities. This algorithm is particularly well-suited for datasets where feature-independence assumptions hold to a reasonable extent. Despite its inherent simplicity and the naive assumption of feature independence, Gaussian NB often performs remarkably well, especially in scenarios with limited training data. A recent review including software implementation and extensions of the method can be found in wickramasinghe2021naive.
A.0.5. Random Forest (RF)
Random forest (RF) is an ensemble learning approach for both classification and regression tasks, known for harnessing the collective strength of multiple decision trees to deliver robust and well-generalized predictions (breiman2001random). It operates by creating bootstrap samples from the training data and randomly selecting features during tree induction. The ensemble’s prediction is aggregated through majority voting (classification) or averaging (regression) to reduce variance and improve model accuracy. This ensemble approach effectively addresses overfitting issues common in single decision trees, while maintaining feature importance measures for understanding the impact of individual features on the prediction. However, choosing the optimal number of trees and tuning hyperparameters remain crucial steps in optimizing RF performance in various applications. Although there are newer techniques for pruning RF (manzali2023random), we use the hyperparameter specifications from lee21.
Appendix B Recommendations for Operationalizing Fairness Search
@@
In hopes of trying to amplify best practices and encourage broader adoption of responsible AI development methodologies, this paper also offers four recommendations based on our findings for public and private sector entities trying to develop their risk management frameworks and LDA searches within the existing AI legal landscape.
B.1. Recommendation 1: Develop an LDA Forward Development Process
@@
Any serious movement towards transitioning to automated decision-making systems in the public and private sectors merits utilizing a rigorous, justifiable, and transparent process in the early model development stage. The foundation of this is predicated on a structured and well-documented LDA search. What this looks like might vary based on institutional capacity to establish what is ”reasonable” for their particular case. For some, an initial horizontal LDA search, surveying across models for fairness gains, may be sufficient, while others may want to consider more comprehensive approaches as used in vertical LDA searches, e.g., tuning, or some combination of both. Regardless, a search should be undertaken as the current disparate impact legal regime expects that should precautions not be taken or deemed insufficient, an organization could then expect to bear costs associated with re-evaluating, re-developing, and replacing the product or service with an LDA (black2024). It is preferable then that rather than looking at fairness assessments as an isolated step in the model development cycle, organizations should aim to have interoperable internal systems in place that can responsibly steward and deploy data across the model lifecycle.
@@
B.2. Recommendation 2: Consider a Developer-first Approach to Documentation and Compliance
State-based AI legislation is demonstrating an emerging interest in AI accountability as primarily a developer-oriented responsibility (colorado2024). With model multiplicity being a widely accepted theory in the AI community, there is almost always a less discriminatory algorithm should one look for it; thus the law is increasingly resting the burden of model harm in the development process if efforts to circumvent potential harm are not proven. Against this context, organizations should consider adjusting their risk management and compliance strategies to focus on the role of their engineering team in the product development process.
@@ B.2. Recommendation 2: Consider a Developer-first Approach to Documentation and Compliance
B.3. Recommendation 3: Ensure Adequate Government Support for LDA Search & Development
To ensure that financial institutions conduct robust searches for LDAs, fairness obligations must be accompanied by clear guidance, concrete examples, and best-practice principles (chien2024). Policymakers should provide comprehensive guidelines in key areas, including appropriate debiasing techniques, recommendations for the proper depth of LDA searches, considerations for LDA viability, and suitable fairness metrics in a variety of different contexts (chien2024).
@@ B.3. Recommendation 3: Ensure Adequate Government Support for LDA Search & Development
B.4. Recommendation 4: Industry Peers Should Benchmark Algorithms and Share Best Practices
@@
Each of these recommendations are independent, yet complementary steps for the at-scale adoption of AI accountability in practice. They assist in translating the normative values responsible AI encourages into systems and processes, establishing ubiquitous, standardized AI fairness across the board. As outlined, developers, institutions, and governments each have role-specific responsibilities to develop systemic and structured expectations on the process, documentation, and standards required to achieve this goal. Auditing and compliance work, as a result, can become less opaque, and the use and development of AI systems can be demystified for increased process efficiency and public trust cultivation.
Appendix C Data Visualizations
C.1. HMDA 2021 Home Ownership Race Share for Approved and Denied Loans
Figure A. 1.
C.2. HMDA 2021 Home Ownership Race Share for Average Loan Amounts
Figure A. 2.
C.3. HMDA 2021 Home Ownership Race Share for Average Interest Rates
Figure A. 3.
Appendix D Features used in Model Development